OpenStudy (anonymous):

what equation is parallel to y=-2x-7

OpenStudy (igreen):

@nicolemosley678 Please post your answer choices.

OpenStudy (igreen):

Graph the equation, and your choices here: https://www.desmos.com/calculator

OpenStudy (anonymous):

y=-2x+7 y=-2x+6 y=-2x-2 y=3x-8

OpenStudy (anonymous):

If two equations are parallel, then they both have the exact same slope. The y-intercept can be different, but the slope has to be the same.

OpenStudy (igreen):

Are you sure those are the choices? Because the first three are parallel..

OpenStudy (anonymous):

In this case, it'd be the second one.

OpenStudy (igreen):

@iPwnBunnies Can ya figure this out?

OpenStudy (ipwnbunnies):

What Emily said.

OpenStudy (igreen):

But how is it the second one?

OpenStudy (ipwnbunnies):

Oh, tru. 2 of those equations are parallel, lelelel. Maybe a mistake in typing.

OpenStudy (igreen):

I don't know.. @nicolemosley678 Are you sure those are your choices?

OpenStudy (shamim):

1,2,3 equations frm ur answer choices r parallel with ur given equation of line

OpenStudy (shamim):

look the coofficient of x frm ur given equation is-2

OpenStudy (shamim):

its slope of ur given equation of line

OpenStudy (shamim):

slope m=-2

OpenStudy (shamim):

do u need more explanation